Consider the following diprotic acid H2Z characterized by the following equilibria:

            H2Z(aq) D H+(aq) + HZ(aq) Ka1 = 2 x 10–5 pKa1 = 4.70

            HZ(aq) D H+(aq) + Z2(aq) Ka2 = 4 x 10–8 pKa2 = 7.40

a. Calculate the pH of a solution of 0.085 M H2Z.

b. Calculate the pH of a solution of 0.085 M NaHZ.

c. Calculate the pH of a solution of 0.085 M Na2Z.

d. Calculate the pH of a solution composed of 0.070 M NaHZ and 0.080 M Na2Z.
Answer #1

@ pH of solution 0.085M HzZ ka, = 24јо ka, = 45o 8 CH+] = Kaic - | 2х6°хоо&S - 1-3 o 38) от 3 pH = -1°C +1 = - Je CJ• 30 3&х- 0.08 Na 12 Kaz= 4x10-8 [17] = v Kg xe = √ 4x108*0,085 CH] = 5-8309410-5 pH = -log Cht = - log (5.8309X16-5) Iph = 4.23437 p